... 4.5. III:6A:7!!! Homework-WAUCM (weighted average unit contribution margin) (CVP-cost volume profit analysis) Last mo
Answer #1

A) 11.33

WAUCM=contribution per sales mix/number of sales mix

number of units sold=total sales/unit price

1. baseball bats=1000/20=50 units

2. tennis rackets=2000/20=100 units

3. bowling balls=3000/20=150 units

so the sales mix is 50:100:150 on simplyfying 1:2:3

number of mix=1+2+3=6

contribution per unit= contribution/number of units sold

contribution=sales-total VC

   1. baseball bats=1000-600=400

2. tennis rackets=2000-1300=700

3. bowling balls=3000-750=2250

contribution per unit

  1. baseball bats=400/50=8

2. tennis rackets=700/100=7

3. bowling balls=2250/150=15

contribution per sales mix=contribution from sale of 1 unit of baseball bats + contribution from sale of 2 unit of tennis racket + contribution from sale of 3 unit of bowling balls

contribution per sales mix=(1*7)+(2*8)+(3*15)=68

WAUCM=68/6=11.33

B) breakeven points in units is 110 units

breakeven units=total fixed cost/WAUCM

breakeven units=1250/11.33=110.32 rounding to nearest whole number 110 units

for the CFIS

sales mix is 1:2:3

break even sales for each product is

1. baseball bats=110*(1/6)=18.33 units rounding to nearest whole number 18 units

2. tennis rackets=110*(2/6)=36.66 units rounding to nearest whole number 37 units

3. bowling balls=110*(3/6)= 55 units

C)

if sports co needs a targeted profit of 3500 then the sales in units would be=(total fixed cost +3500)/WAUCM

=(1250+3500)/11.33=419.24 units rounding to nearest whole number 419 units

for CFIS

1. baseball bats=419*(1/6)=69.83 units rounding to nearest whole number 70 units

2. tennis rackets=419*(2/6)=139.66 units rounding to nearest whole number 140 units

3. bowling balls=419*(3/6)= 209.5 units rounding to nearest whole number 209 units

in sales= break even units*selling price

1. baseball bats= 70*20=1400

2. tennis rackets= 140*20=2800

3. bowling balls= 209*20=4180

total breakeven sales=4180+2800+1400=8380

D) last month profit is 2100

to determine last month's profit or loss=total sales-total VC-total fixed cost

total sales=1000+2000+3000=6000

total VC=600+1300+750=2650

total fixed cost=1250

last month's profit=6000-2650-1250=2100